OpenStudy (anonymous):

Martina runs 3 miles in 28 minutes. At the same rate, how many miles would she run in 42 minutes

OpenStudy (anonymous):

\[\frac{ 3 miles }{ 28 minutes} = \frac{ ? miles }{ 42 minutes }\] Cross multiply to get 42*3 = 28*x Divide both sides by 28 to get 42/28*3 = x, or x = 4.5 miles

OpenStudy (anonymous):

thank you very much